OpenStudy (anonymous):

how would i solve these radicals: sqrt of 6 multiplied by the sqrt of 26

OpenStudy (anonymous):



OpenStudy (btaylor):

\[\sqrt{6} \times \sqrt{26} = \sqrt{6 \times 26} = \sqrt{3 \times 2 \times 2 \times 13} = 2 \sqrt{6 \times 13} = 2\sqrt{78} \approx 12.49\]